OpenStudy (anonymous):

try and solve this i bet u cant! 8 and 1\5 - 3 and 1\4

OpenStudy (anonymous):

i can...

OpenStudy (anonymous):

you lose the bet...:)

OpenStudy (eyust707):

hehe

OpenStudy (anonymous):

i keep getting 5 1\20

OpenStudy (eyust707):

I bet that YOU can!

OpenStudy (anonymous):

keep trying... i bet you can do this too...

OpenStudy (anonymous):

try common denominators...

OpenStudy (anonymous):

I can't, so you won the bet.

OpenStudy (anonymous):

20 is a common denominator

OpenStudy (anonymous):

ok so that 1/5 becomes ??/20 and the 1/4 becomes ???/20

OpenStudy (anonymous):

\[\frac{41}{5}-\frac{13}{4}=4\frac{19}{20} \]

OpenStudy (anonymous):